OpenStudy (anonymous):

Suppose you are saving money for a new cell phone. At week 0 you have $20 and at week 10 you have $100. If y represents the amount of money you have saved and x represents the number of weeks, find and interpret the slope of the linear equation. HELP , pleaseee? 1/8; $8 is put away each week. 10; $10 is put away each week. 8; $8 is put away each week. 1/10; $10 is put away each week.

hero (hero):

The 1st point is (x,y) = (0,20) The 2nd point is (x,y) = (10,100) Which means now you can simply use y = mx + b to solve: 20 = 0m + b 100 = 10m + b Isolating and solving for b you get: b = 20 b = 100 - 10m Next set b = b 100 - 10m = 20 Afterwards isolate and solve for m to get: 100 - 20 = 10m 80 = 10m 80/10 = m 8/1 = m in this case, y/x = m means money/week

hero (hero):

So what is the answer @andreascott

OpenStudy (anonymous):

well 8 divided by 1 is 8 . So, its the third one? Right ?